An Iron Rod Part A On a part ime job you are asked to bring a cyindical ron rod of densly 7800k/gth 91.1 can and dameter 2 30 c due to gravity, g 981 m/s Express your answer numerically in newtons from a storage r room to a machinist Cakulate the weight of the red w The acceleration
Answer #1

We know that mass of the cylindrical rod

Mass=density × Volume

M=dV. ....1 Here V=πr^2 ×h

here d=7800 kg /m^3. h =91.1cm =0.911m, r=2.3/2cm

r=0.0115 m

so V=π ×(.0115)^2 ×0.911 =3.785×10^(-4) m^3

from equation 1

M=7800×3.785×10^(-4)=2.9523 kg

M=2.95 kg

weight of the cylinder= w= Mg

w=2.95×9.81=28.9395 N

w=28.94 N. Answer